# Homeomorphisms of continua through projective Fraïssé limits

### v1

@@ 1. $\{(x,x)\mid x\in X\}\subseteq U$ for each $U\in{\mathcal{U}}$;
(2.5) $\forall x,y\in\mathbb{K}\;\big(\phi(x)R\phi(y)\Rightarrow d^{0}(\pi(x),\pi(y))<\epsilon/2\big).$
@@
3. Joint Projection Property
Let $\mathcal{K}$ be a category. Following Kechris–Rosendal [8, Section 2], and adapting their work to the projective setting, we define the category ${\mathcal{K}}_{\textrm{p}}$ as follows.
@@ 3. Joint Projection Property
if for each $i$ the map $\alpha$ is an epimorphism (an approximate epimorphism, respectively) between $(A,B,f_{i},g_{i})$ and $(A^{\prime},B^{\prime},f^{\prime}_{i},g^{\prime}_{i})$.
@@
Suppose that $\mathcal{K}$} is a transitive projective Fraïssé class with the property that $\rm Aut({\mathbb{K}})$} has a dense image under the canonical homomorphism into ${\rm Homeo}(K)$. Let $n\in{\mathbb{N}}$.
@@ 3. $(iii)$
Note that $(iii)$ is the dualized version of Kechris-Rosendal.
@@
We turn to $(ii)$. For the non-empty open set $U\subseteq{\rm Homeo}(K)^{n}$ there again exist $(A^{\prime},B^{\prime},f^{\prime}_{i},g^{\prime}_{i})\in{\mathcal{K}}_{\textrm{p}}$, $i<n$, and $\varphi^{\prime}\in{\rm Epi}({\mathbb{K}},A^{\prime})$ with
@@
Fix a non-empty open set $V\subseteq{\rm Aut}({\mathbb{K}})$. By Corollary 2.5. and Lemma 3.7, we can assume that for some $(A,B,f_{i},g_{i})\in{\mathcal{K}}_{\textrm{p}}$, $i<n$, and $\varphi\in{\rm Epi}({\mathbb{K}},A)$, we have
@@
By the half-approximate JPP, we obtain $A^{+},B^{+}\in{\mathcal{K}}$, $f^{+}_{i},g^{+}_{i}\in{\rm Epi}(A^{+},B^{+})$, $i<n$, and $\mathcal{K}$}-epimorphisms $\alpha:A^{+}\to A$, $\alpha^{\prime}:A^{+}\to A^{\prime}$, such that, for each $i<n$, $\alpha$ is a ${\mathcal{K}}_{\textrm{p}}$-epimorphism from $(A^{+},B^{+},f_{i}^{+},g_{i}^{+})$ to $(A,B,f_{i},g_{i})$ and $\alpha^{\prime}$ is an approximate ${\mathcal{K}}_{\textrm{p}}$-epimorphism to $(A^{\prime},B^{\prime},f^{\prime}_{i},g^{\prime}_{i})$.
@@
Suppose that $\mathcal{K}$} is a transitive projective Fraïssé class with the property that $\rm Aut({\mathbb{K}})$} has a dense image under the canonical homomorphism into ${\rm Homeo}(K)$, and let $n\in{\mathbb{N}}$.
@@
By our assumptions there are $\varsigma\in U$, $\varsigma^{\prime}\in U^{\prime}$ that are conjugate to each other, i.e. for some $\tau_{0}\in{\rm Homeo}(K)$
@@
and $\varphi$, $\varphi^{\prime}\circ\tau$ factor through $g^{+}\circ\varphi^{+}$ (therefore they factor through $\varphi^{+}$, too). This can be done by picking $\psi^{*}\in{\rm Epi}({\mathbb{K}},C)$ such that $(\varphi\times\varphi^{\prime}\circ\tau)$ factors through $\psi^{*}$, and applying Lemma 3.8 to $\varphi^{*}=\psi^{*}\circ\sigma$ and $\psi^{*}$. If $f^{+}\circ\varphi^{+}=\psi^{*}\circ\sigma^{-1}$ and $g^{+}\circ\varphi^{+}=\psi^{*}$, then as $\varphi$, $\varphi^{\prime}\circ\tau$ factor through $\psi^{*}$, i.e. $\varphi=\delta\circ\psi^{*}$, $\varphi^{\prime}\circ\tau=\delta^{\prime}\circ\psi^{*}$ for some $\delta$ and $\delta^{\prime}$, clearly
@@
so we need to check this for all $x$, i.e.
@@
so we need to verify that this holds for all $x$, i.e.
@@ 4. Dense conjugacy classes in products—the pseudoarc case
We refer the reader to Section 1 for the definition and relevant properties of $\mathcal{P}$}—the category of finite reflexive linear graphs. By $\mathbb{P}$} we denote the projective Fraïssé limit of $\mathcal{P}$}. In this case, the canonical continuous homomorphism ${\rm pr}:{\rm Aut}({\mathbb{P}})\to{\rm Homeo}(P)$ is injective, and we will identify ${\rm Aut}({\mathbb{P}})$ with its image, that is,
@@
  • the map $\alpha$ is an epimorphism between $(A^{+},B^{+},f_{i}^{+},g_{i}^{+})$ and $(A,B,f_{i},g_{i})$,
  • the map $\alpha^{\prime}$ is an approximate epimorphism between $(A^{+},B^{+},f_{i}^{+},g_{i}^{+})$ and $(A^{\prime},B^{\prime},f^{\prime}_{i},g^{\prime}_{i})$.
@@ 1. $(enumi)_{1}$
    $f^{-1}(b)$ is the union of intervals each is of length at least $2$,
since exact epimorphisms are closed under composition. We introduce the notation $\mathbf{R}$$(.,.)$ for division with reminder, and write $\mathbf{R}(j,m)=k$ iff $k\in[0,m)$ and $j\equiv k$ mod $m$.
First we define $A^{+}$ and $B^{+}$, we start with the latter. Let $m=|B|$, $m^{\prime}=|B^{\prime}|$, that is, $B$ is a path of length $m$, $B^{\prime}$ is a path of length $m^{\prime}$. We set $B^{+}$ to be a path of length $mm^{\prime}$. Identifying $B^{\prime}$ with $0,m^{\prime})$ and $B^{+}$ with $0,mm^{\prime})$, define $\beta^{\prime}:B^{+}\to B^{\prime}$ by the equation
  1. $(enumi)_{2}$
    $\beta^{\prime}(j)=\lfloor j/m\rfloor$.
Next we define $\beta:B^{+}\to B$, which will satisfy that whenever $\mathbf{R}(j,m)=k$, then
$\beta(j)\in\{k,m-1-k\}.$
The epimorphism $\beta$ is given by
  1. $(enumi)_{3}$
    $\beta(j)=\mathbf{R}(j,m)$,   for $j\in[2lm,(2l+1)m)$ with $0\leq 2l\leq m^{\prime}-1$;
  2. $(enumi)_{4}$
    $\beta(j)=m-1-\mathbf{R}(j,m)$,   for $j\in[(2l+1)m,(2l+2)m)$ with $0\leq 2l\leq m^{\prime}-2$.
The reflexive linear graph $A^{+}$ and the epimorphisms $\alpha$, $\alpha^{\prime}$ are defined similarly. If $n=|A|$, $n^{\prime}=|A^{\prime}|$, then let $A^{+}$ to be a path of length $nn^{\prime}$. Define $\alpha^{\prime}:A^{+}\to A^{\prime}$ by the equation
  1. $(enumi)_{5}$
    $\alpha^{\prime}(j)=\lfloor j/n\rfloor$.
The map $\alpha:A^{+}\to A$ is defined by
  1. $(enumi)_{6}$
    $\alpha(j)=\mathbf{R}(j,n)$,  for $j\in[2ln,(2l+1)n)$ with $0\leq 2l\leq n^{\prime}-1$,
  2. $(enumi)_{7}$
    $\alpha(j)=n-1-\mathbf{R}(j,n)$,  for $j\in[(2l+1)n,(2l+2)n)$ with $0\leq 2l\leq n^{\prime}-2$.
@@ 2. $(enumi)_{7}$
(4.1) $\beta\circ f^{+}\$$=\ f\circ\alpha,$
(4.2) $\beta^{\prime}\circ f^{+}\$$R\ f^{\prime}\circ\alpha^{\prime}.$
We note that if $J\subseteq A^{+}$ is an interval satisfying $f^{+}[J]\subseteq[lm,(l+1)m)$, for some $l<m^{\prime}$, then $f^{+}{\restriction}J$ is determined by $f$ (by the requirement (4.1) and by injectivity of $\beta$ on $[lm,(l+1)m)$ guaranteed by $(enumi)_{3}$, $(enumi)_{4}$).
We will define the sequence $s_{0},s_{1},\ldots,s_{n^{\prime}-1}$ so that
  1. $(I)$
    $s_{j}\in[jn,(j+1)n-1)$,
  2. $(II)$
    for $a<n^{\prime}n$,
    $\beta\circ f^{+}(a)=f\circ\alpha(a),$
  3. $(III)$
    if $j<n^{\prime}-1$, then, for $a\in[jn,s_{j}]$,
    $\beta^{\prime}\circ f^{+}(a)=f^{\prime}\circ\alpha^{\prime}(a),$
    and for $a\in[s_{j}+1,(j+1)n)$,
    $\beta^{\prime}\circ f^{+}(a)=f^{\prime}(\alpha^{\prime}(a)+1),$
    moreover, for $j=n^{\prime}-1$ and $a\in[jn,(j+1)n)$,
    $\beta^{\prime}\circ f^{+}(a)=f^{\prime}\circ\alpha^{\prime}(a).$
We note that $(II)$, $(III)$ imply (4.2), since $f^{\prime}(\alpha^{\prime}(a))Rf^{\prime}(\alpha^{\prime}(a)+1)$ because $f^{\prime}$ is an epimorphism. So it remains to prove the following statement.
Suppose that $j=0$, or $j>0$ and $s_{0},s_{1},\ldots,s_{j-1}$, $f^{+}{\restriction}[0,jn)$ are defined so that $(II)$, $(III)$ hold. Then it is possible to define $s_{j}$ and $f^{+}{\restriction}[0,(j+1)n)$ so that they similarly satisfy $(II)$, $(III)$.
The statement is proved by induction on $j$. First we pick $f^{+}(jn)\in B^{+}$.
If $j=0$, then we use that $\beta\times\beta^{\prime}$ is a bijection between $B^{+}$ and $B\times B^{\prime}$ and let $f^{+}(0)$ be the unique value in $B^{+}$ that satisfies
[…]
If $j>0$, then by ($(III)$) we have
$\beta^{\prime}\circ f^{+}(jn-1)=f^{\prime}\circ(\alpha^{\prime}(jn-1)+1),$
and by the definition of $\alpha^{\prime}$ we have $\alpha^{\prime}(jn-1)+1=j-1+1=\alpha^{\prime}(jn)$. Moreover, the way we defined $\alpha$ implies $\alpha(jn-1)=\alpha(jn)$, therefore letting $f^{+}(jn)=f^{+}(jn-1)$ will ensure (4.1) and $(II)$. If $j<n^{\prime}-1$, then consider the value
$c=f^{\prime}(\alpha^{\prime}(jn)+1)=f^{\prime}(\alpha^{\prime}((j+1)n))=f^{\prime}(j+1).$
If $c=f^{\prime}(\alpha^{\prime}(jn))$, or $j=n^{\prime}-1$ and $c$ is not defined then we just define $f^{+}{\restriction}[jn,(j+1)n)$ so that $\beta^{\prime}\circ f^{+}{\restriction}[jn,(j+1)n)=c$, and $\beta\circ f^{+}{\restriction}[jn,(j+1)n)=f\circ\alpha{\restriction}[jn,(j+1)n)$. (In this case $s_{j}$ can be just $jn$.)
Otherwise, $j<n^{\prime}-1$ and $c\in\{f^{\prime}(\alpha^{\prime}(jn))-1,f^{\prime}(\alpha^{\prime}(jn))+1\}$ hold necessarily, and our aim is to ultimately get $\beta^{\prime}(f^{+}((j+1)n-1))=c$. For simplicity we assume that
(4.3) $c=f^{\prime}(\alpha^{\prime}(jn))+1$
@@ 3. $(III)$
(4.4) $f^{\prime}(\alpha^{\prime}(jn))(=\beta^{\prime}(f^{+}(jn)))\ {\rm is\ even,}$
the other three cases are handled the exact same way. By these assumptions and by the way $\beta^{\prime}$ was defined ($(enumi)_{2}$) we note that
$(\beta^{\prime})^{-1}(f^{\prime}(\alpha^{\prime}(jn))=[f^{\prime}(\alpha^{\prime}(jn)n,(f^{\prime}(\alpha^{\prime}(jn))+1)n),$
and
$(\beta^{\prime})^{-1}(f^{\prime}(\alpha^{\prime}(jn)+1)=[(f^{\prime}(\alpha^{\prime}(jn)+1)n,(f^{\prime}(\alpha^{\prime}(jn))+2)n).$
Now we let $s_{j}\geq jn$ be minimal such that $f(\alpha(s_{j}))=m-1$, and define $f^{+}{\restriction}[jn,s_{j}]$ so that $\beta^{\prime}\circ f^{+}{\restriction}[jn,s_{j}]\equiv f^{\prime}(\alpha^{\prime}(jn))$, and $\beta\circ f^{+}{\restriction}[jn,s_{j}]=f\circ\alpha{\restriction}[jn,s_{j}]$. It follows that $f^{+}(s_{j})=f^{\prime}(\alpha^{\prime}(jn)+1)n-1$. Using the assumption $(enumi)_{1}$ on $f$ we obtain that $s_{j}+1<(j+1)n$. Let $f^{+}(s_{j}+1)=f^{+}(s_{j})+1$, using $(enumi)_{3}$
$\beta(f^{+}(s_{j}+1))=\beta(f^{+}(s_{j})+1)=\beta(f^{\prime}(\alpha^{\prime}(jn)+1)n-1+1)=\beta(f^{\prime}(\alpha^{\prime}(jn)+1)n),$
and note that $f^{\prime}(\alpha^{\prime}(jn)+1)=c=f^{\prime}(\alpha^{\prime}(jn))+1$ is odd by our assumption (4.4), therefore (by $(enumi)_{4}$)
$\beta(f^{+}(s_{j}+1))=\beta(f^{\prime}(\alpha^{\prime}(jn)+1)n)=m-1,$
so $\beta(f^{+}(s_{j}+1))=f(\alpha(s_{j}))$. Using that $s_{j}\geq jn$ is minimal with the condition $f(\alpha(s_{j}))=m-1$ and recalling $(enumi)_{1}$, $(enumi)_{6}$ we get $s_{j}<(j+1)n-1$,
$f(\alpha(s_{j}+1))=f(\alpha(s_{j})+1)=f(\alpha(s_{j}))=m-1,$
hence $(II)$ remains true. It is easy to check that $\alpha^{\prime}(s_{j})=\alpha^{\prime}(s_{j}+1)$ since $jn+1\leq s_{j}+1<(j+1)n$ and so
$\beta^{\prime}(f^{+}(s_{j}+1))=\beta^{\prime}(f^{+}(s_{j}))+1=f^{\prime}(\alpha^{\prime}(s_{j}))+1=f^{\prime}(\alpha^{\prime}(s_{j}+1))+1,$
and we get ($(III)$).
Finally, we define $f^{+}{\restriction}[s_{j}+1,(j+1)n)$ so that $\beta^{\prime}\circ f^{+}{\restriction}[s_{j}+1,(j+1)n)\equiv\beta^{\prime}(f^{+}(s_{j}+1))$, and $\beta\circ f^{+}{\restriction}[s_{j}+1,(j+1)n)=f\circ\alpha{\restriction}[s_{j}+1,(j+1)n)$.
The statement is proved and the lemma follows. ∎

### v2

@@ 2. Topologies on ${\rm Aut}({\mathbb{K}})$
For notion not defined in this section, the reader should refer to Appendix A.
@@ 1. $\{(x,x)\mid x\in X\}\subseteq U$ for each $U\in{\mathcal{U}}$;
(2.5) $\forall x,y\in\mathbb{K}\;\big(\phi(x)R\phi(y)\Rightarrow d^{0}({\rm pr}(x),{\rm pr}(y))<\epsilon/2\big).$
@@
3. Joint Projection Properties
3(A). The formulation of Joint Projection Properties
Let $\mathcal{K}$ be a category. Following Kechris–Rosendal [10, Section 2], and adapting their work to the projective setting, we define the category ${\mathcal{K}}_{\textrm{p}}$ as follows.
@@ 3. Joint Projection Properties
if for each $i<n$ the map $\alpha$ is an epimorphism (an approximate epimorphism, respectively) between $(A,B,f_{i},g_{i})$ and $(A^{\prime},B^{\prime},f^{\prime}_{i},g^{\prime}_{i})$.
@@
3(B). Joint Projection Properties and density of orbits for diagonal conjugacy actions
@@
Suppose that $\mathcal{K}$} is a transitive projective Fraïssé class with the property that the image of $\rm Aut({\mathbb{K}})$} under the canonical homomorphism $\rm pr$ is dense in ${\rm Homeo}(K)$. Let $n\in{\mathbb{N}}$.
@@ 3. $(iii)$
Note that $(iii)$ is the dualized version of [10].
@@
We turn to $(ii)$. Fix a non-empty open set $U\subseteq{\rm Homeo}(K)^{n}$. We can again assume that there exist $(A^{\prime},B^{\prime},f^{\prime}_{i},g^{\prime}_{i})\in{\mathcal{K}}_{\textrm{p}}$, $i<n$, and $\varphi^{\prime}\in{\rm Epi}({\mathbb{K}},A^{\prime})$ with
@@
Fix a non-empty open set $V\subseteq{\rm Aut}({\mathbb{K}})$. By Corollary 2.5 and Lemma 3.8, we can assume that for some $(A,B,f_{i},g_{i})\in{\mathcal{K}}_{\textrm{p}}$, $i<n$, and $\varphi\in{\rm Epi}({\mathbb{K}},A)$, we have
@@
By the half-approximate JPP, we obtain $A^{+},B^{+}\in{\mathcal{K}}$, $f^{+}_{i},g^{+}_{i}\in{\rm Epi}(A^{+},B^{+})$, $i<n$, and $\mathcal{K}$}-epimorphisms $\alpha:A^{+}\to A$, $\alpha^{\prime}:A^{+}\to A^{\prime}$ such that, for each $i<n$, $\alpha$ is a ${\mathcal{K}}_{\textrm{p}}$-epimorphism from $(A^{+},B^{+},f_{i}^{+},g_{i}^{+})$ to $(A,B,f_{i},g_{i})$ and $\alpha^{\prime}$ is an approximate ${\mathcal{K}}_{\textrm{p}}$-epimorphism to $(A^{\prime},B^{\prime},f^{\prime}_{i},g^{\prime}_{i})$.
@@
Suppose that $\mathcal{K}$} is a transitive projective Fraïssé class with the property that the image of $\rm Aut({\mathbb{K}})$} under the canonical homomorphism $\rm pr$ is dense in ${\rm Homeo}(K)$. Let $n\in{\mathbb{N}}$.
@@
By our assumptions there are $\varsigma\in U$, $\varsigma^{\prime}\in U^{\prime}$ that are conjugate to each other, that is, for some $\tau_{0}\in{\rm Homeo}(K)$
@@
and $\varphi$, $\varphi^{\prime}\circ\tau$ factor through $g^{+}\circ\varphi^{+}$ (therefore they factor through $\varphi^{+}$, too). This can be done by picking $\psi^{*}\in{\rm Epi}({\mathbb{K}},C)$ such that $(\varphi\times\varphi^{\prime}\circ\tau)$ factors through $\psi^{*}$, and applying Lemma 3.9 to $\varphi^{*}=\psi^{*}\circ\sigma$ and $\psi^{*}$. If $f^{+}\circ\varphi^{+}=\psi^{*}\circ\sigma^{-1}$ and $g^{+}\circ\varphi^{+}=\psi^{*}$, then as $\varphi$, $\varphi^{\prime}\circ\tau$ factor through $\psi^{*}$, that is, $\varphi=\delta\circ\psi^{*}$, $\varphi^{\prime}\circ\tau=\delta^{\prime}\circ\psi^{*}$ for some $\delta$ and $\delta^{\prime}$, clearly
@@
so we need to check this for all $x$, that is,
@@
so we need to verify that this holds for all $x$, that is,
@@
3(C). Comeager conjugacy class in ${\rm Homeo}(K)$
The following definition and theorem are the dualized versions of the corresponding notions in [10].
Definition 3.16.
Let $\mathcal{K}$} be a projective Fraïssé class. We say that ${\mathcal{K}}_{\rm p}$ satisfies the Weak Projective Amalgamation Property (in short, WPAP) if for each $(A,B,f,g)\in{\mathcal{K}}_{\rm p}$ there exist $(A^{+},B^{+},f^{+},g^{+})\in{\mathcal{K}}_{\rm p}$ and an epimorphism
$\alpha:\ (A^{+},B^{+},f^{+},g^{+})\to(A,B,f,g)$
such that for every $(A^{\prime},B^{\prime},f^{\prime},g^{\prime}),(A^{\prime\prime},B^{\prime\prime},f^{\prime\prime},g^{\prime\prime})\in{\mathcal{K}}$ and epimorphisms
[…]
there exist $(A^{*},B^{*},f^{*},g^{*})\in{\mathcal{K}}$ and epimorphisms
[…]
such that
$\alpha\circ\beta^{\prime}\circ\gamma^{\prime}=\alpha\circ\beta^{\prime\prime}\circ\gamma^{\prime\prime}\ \textrm{(as mappings in }{\rm Epi}(A^{*},A){\rm)}.$
The interested reader may dualize the proof [10, Theorem 3.4] to obtain the following.
Theorem 3.17.
Suppose that $\mathcal{K}$} is a projective Fraïssé class, let $\mathbb{K}$} be its limit. Then the following are equivalent
  1. $(i)$
    There exists a comeager conjugacy class in ${\rm Aut}({\mathbb{K}})$.
  2. $(ii)$
    ${\mathcal{K}}_{\rm p}$ has the JPP and WPAP.
We remark that the above equivalence can be adapted to the approximate setting to give an equivalent condition for the existence of the generic homeomorphism.
Definition 3.18.
Let $\mathcal{K}$} be a projective Fraïssé class. We say that ${\mathcal{K}}_{\rm p}$ satisfies the approximate WPAP if for each $(A,B,f,g)\in{\mathcal{K}}_{\rm p}$ there exist $(A^{+},B^{+},f^{+},g^{+})\in{\mathcal{K}}_{\rm p}$ and an approximate epimorphism
$\alpha:\ (A^{+},B^{+},f^{+},g^{+})\to(A,B,f,g)$
such that for every $(A^{\prime},B^{\prime},f^{\prime},g^{\prime}),(A^{\prime\prime},B^{\prime\prime},f^{\prime\prime},g^{\prime\prime})\in{\mathcal{K}}$ and approximate epimorphisms
[…]
there exist $(A^{*},B^{*},f^{*},g^{*})\in{\mathcal{K}}$ and approximate epimorphisms
[…]
such that
$(\alpha\circ\beta^{\prime}\circ\gamma^{\prime})\ R\ (\alpha\circ\beta^{\prime\prime}\circ\gamma^{\prime\prime})\ \textrm{(in }{\rm Epi}(A^{*},A){\rm)}.$
Using the proof of [10, Theorem 3.4] coupled with our ideas it is possible to show the following.
Theorem 3.19.
Suppose that $\mathcal{K}$} is a transitive projective Fraïssé class, let $K={\mathbb{K}}/R^{\mathbb{K}}$ be the quotient compactum, and assume that ${\rm Aut}({\mathbb{K}})$ is dense in ${\rm Homeo}(K)$. Then the following are equivalent
  1. $(i)$
    There exists a comeager conjugacy class in ${\rm Homeo}(K)$.
  2. $(ii)$
    ${\mathcal{K}}_{\rm p}$ has the approximate JPP and the approximate WPAP.
Remark 3.20.
To make the above definition meaningful, we point out two facts in regard to composing approximate epimorphisms. First, it is not difficult to see that Remark 3.7 implies that if $\delta$ is the composition of two approximate epimorphism between $(A^{\prime},B^{\prime},f^{\prime},g^{\prime})$ and $(A,B,f,g)$, then $f^{\prime}(a_{0})=g^{\prime}(a_{1})$ implies $f(\delta(a_{0}))R^{4}g(\delta(a_{1}))$.
Second, in the above theorem one can also require that the approximate epimorphism $\alpha:\ (A^{+},B^{+},f^{+},g^{+})\to(A,B,f,g)$ satisfy the stronger condition
$(f^{+}(a_{0})R^{4}g^{+}(a_{1}))\ \Rightarrow\ (f(\alpha(a_{0}))Rg(\alpha(a_{1})),$
therefore $(\alpha\circ\beta^{\prime}\circ\gamma^{\prime})$ and $(\alpha\circ\beta^{\prime\prime}\circ\gamma^{\prime\prime})$ are approximate epimorphisms.
@@ 4. Dense conjugacy classes in products—the pseudoarc case
We refer the reader to Section 1 for the definition and relevant properties of $\mathcal{P}$}—the category of finite reflexive linear graphs. By $\mathbb{P}$} we denote the projective Fraïssé limit of $\mathcal{P}$}. In this case, the canonical continuous homomorphism ${\rm pr}:{\rm Aut}({\mathbb{P}})\to{\rm Homeo}(P)$ is injective as the class $\mathcal{P}$} obviously fulfills condition (iv) of Proposition A.2; so it fulfills the conclusions of Proposition A.5. We identify ${\rm Aut}({\mathbb{P}})$ with its image, that is,
@@
  • the map $\alpha$ is an approximate epimorphism between $(A^{+},B^{+},f_{i}^{+},g_{i}^{+})$ and $(A,B,f_{i},g_{i})$,
  • the map $\alpha^{\prime}$ is an epimorphism between $(A^{+},B^{+},f_{i}^{+},g_{i}^{+})$ and $(A^{\prime},B^{\prime},f^{\prime}_{i},g^{\prime}_{i})$.
@@ 1. $(enumi)_{1}$
    $(f^{\prime})^{-1}(b)$ is the union of intervals each is of length at least $2$,
since exact epimorphisms are closed under composition.
Let
$A=\{0,\dots,m\},\;A^{\prime}=\{0,\dots,m^{\prime}\},\;B=\{0,\dots,n\},\;B^{\prime}=\{0,\dots,n^{\prime}\},$
we define $A^{+}$ with $R^{A^{+}}$. Let
[…]
and $A^{+}=\bigcup_{i\in A}A^{\prime}_{i}$. We define $R^{A^{+}}$ on these sets to be the smallest reflexive and symmetric relation with
$(i,j)\,R^{A^{+}}(i,j^{\prime})\Leftrightarrow|j-j^{\prime}|\leq 1,\;\hbox{ for }i\in A,$
(so that with these definitions each $A_{i}^{\prime}$ is a reflexive linear graph with $m^{\prime}+1$ many elements), and
$(i,m^{\prime})\,R^{A^{+}}(i+1,m^{\prime}),\;\hbox{ for }i<m,\ i\hbox{ is even}.$
$(i,0)\,R^{A^{+}}(i+1,0),\;\hbox{ for }i<m,\ i\hbox{ is odd}.$
Similarly, we let
[…]
and $B^{+}=\bigcup_{i\in B}B^{\prime}_{i}$, where we stipulate that $R^{B^{+}}$ satisfy
$(k,l)\,R^{B^{+}}(k,l^{\prime})\Leftrightarrow|l-l^{\prime}|\leq 1,\;\hbox{ for }k\in B,$
and
$(k,n^{\prime})\,R^{B^{+}}(k+1,n^{\prime}),\;\hbox{ for }k<n,\ k\hbox{ is even},$
$(k,0)\,R^{B^{+}}(k+1,0),\;\hbox{ for }k<n,\ k\hbox{ is odd}.$
Now we let $\alpha\colon A^{+}\to A$ and $\alpha^{\prime}\colon A^{+}\to A^{\prime}$ be defined by
$\alpha(i,j)=i\;\hbox{ and }\;\alpha^{\prime}(i,j)=j.$
Similarly, $\beta\colon B^{+}\to B$ and $\beta^{\prime}\colon B^{+}\to B^{\prime}$ is defined by
$\beta(k,l)=k\;\hbox{ and }\;\beta^{\prime}(k,l)=l.$
It is immediate that $\alpha,\alpha^{\prime},\beta,\beta^{\prime}$ are epimorphisms.
@@
(4.1) $\beta\circ f^{+}\$$R\ \ f\circ\alpha,$
(4.2) $\beta^{\prime}\circ f^{+}\$$=\ f^{\prime}\circ\alpha^{\prime}.$
Let $a^{*}_{0}=(0,0)$, and for $0<i\in A$ define $a^{*}_{i}\in A^{\prime}_{i}$ be the element adjacent to $A_{i-1}^{\prime}$. We define $f^{+}$ by induction so that for each $i$
(4.3) $\beta\circ f^{+}(a^{*}_{i})\$$=\ f\circ\alpha(a_{i}^{*}),$
(4.4) $\beta\circ f^{+}(a)\$$R\ \ \beta\circ f^{+}(a_{i}^{*}),\ (a\in A_{i}^{\prime})$
(4.5) $\beta^{\prime}\circ f^{+}{\restriction}A_{i}^{\prime}\$$=\ f^{\prime}\circ\alpha^{\prime}{\restriction}A_{i}^{\prime}.$
Observe that $f\circ\alpha{\restriction}A_{i}^{\prime}$ is constant, since $\alpha{\restriction}A_{i}^{\prime}$ is constant for each $i\in A$, therefore (4.3) and (4.4) together imply (4.1).
Recall that $a_{0}^{*}=(0,0)$, let
$f^{+}(a_{0}^{*})=(f(0),f^{\prime}(0)),$
so clearly $\beta(f^{+}(a_{0}^{*}))=f(0)=f(\alpha(a_{0}^{*}))$, similarly with $\beta^{\prime}$, $\alpha^{\prime}$.
Suppose that $f^{+}{\restriction}A_{0}^{\prime}\cup\ldots\cup A_{i-1}^{\prime}\cup\{a_{i}^{*}\}$ has been defined with the relevant parts of (4.3)-(4.5) being satisfied.
Let $\varepsilon=f(i+1)-f(i)\in\{-1,0,1\}$. If $\varepsilon=0$, then $f\circ\alpha{\restriction}A^{\prime}_{i}\cup\{a_{i+1}^{*}\}\equiv f(i)=f(i+1)$, and we can define
$f^{+}(a)=(f(i),f^{\prime}(\alpha^{\prime}(a)))\hbox{ for }a\in A^{\prime}_{i}\cup\{a_{i+1}^{*}\}.$
Otherwise, if $\varepsilon\neq 0$, then we distinguish eight cases depending on the parity of $f(i)$ and $i$. Suppose for simplicity that $\varepsilon=1$ (so $f(i+1))=f(i)+1$) and $f(i)$, $i$ are even.
By our hypothesis, $f^{+}(a_{i}^{*})\in B^{\prime}_{f(i)}$, therefore it suffices to define $f^{+}{\restriction}A_{i}\cup\{a_{i+1}^{*}\}$ so that
(4.6) $\begin{array}[]{rl}f^{+}[A_{i}\cup\{a_{i+1}^{*}\}]&\subseteq B^{\prime}_{f(i)}\cup B^{\prime}_{f(i)+1},\\ f^{+}(a_{i+1}^{*})&\in B^{\prime}_{f(i)+1}\end{array}$
alongside the condition (4.5). (This is because $\beta{\restriction}B^{\prime}_{k}\equiv k$). Since $f(i)$ is even we note that
(4.7) $(f(i),n^{\prime})\ R^{B^{+}}\ (f(i)+1,n^{\prime}).$
Now $a_{i}^{*}$ is the endpoint of the finite linear graph $A_{i}^{\prime}=\{(i,j):\ j\in A^{\prime}\}$ that is connected with an element of $A^{\prime}_{i-1}$, and since $i$ is even, we have $a_{i}^{*}=(i,0)$. Let $j_{0}\in A^{\prime}$ be the smallest such that $f^{\prime}(j_{0})=n^{\prime}$, so that by $(enumi)_{1}$
$f^{\prime}(j_{0}+1)=n^{\prime}.$
Letting
$f^{+}(i,j)=\begin{cases}(f(i),f^{\prime}(j)),&\hbox{ for }\;j\leq j_{0};\\ (f(i)+1,f^{\prime}(j)),&\hbox{ for }\;j>j_{0}.\end{cases}$
@@
$f^{+}(a_{i+1}^{*})=f^{+}(i+1,m^{\prime})=(f(i)+1,f^{\prime}(m^{\prime}))\in B^{\prime}_{f(i)+1}$
so by the choice of $j_{0}$ (and (4.7)) we get an epimorphism and (4.3) for $a_{i+1}^{*}$. It is also clear that $f^{+}(i,j),f^{+}(i+1,m^{\prime})\in B^{\prime}_{f(i)}\cup B^{\prime}_{f(i)+1}$ (implying (4.6), so that (4.4) follows). Moreover, $\beta^{\prime}(f^{+}(i,j))=f^{\prime}(j)=f^{\prime}(\alpha^{\prime}(i,j))$ for $j\in A^{\prime}$ and $\beta^{\prime}(f^{+}(i+1,m^{\prime}))=f^{\prime}(m^{\prime})=f^{\prime}(\alpha^{\prime}(i+1,m^{\prime}))$, which implies (4.5).